# MedAide: Towards an Omni Medical Aide via Specialized LLM-based Multi-Agent Collaboration

### v2

@@ 3 Methodology
Figure 1 presents the overall procedure of the proposed MedAide. Here, we detail the systematic workflow procedure, including query rewriting, intent recognition, and agent collaboration.
Figure 1: Illustration of the proposed MedAide framework. MedAide consists of three major phases: (i) query rewriting aims to reformulate medical queries and mine user intents to ensure the system accurately understands the input requirement goals. Among the optimized queries, (ii) intent recognition utilizes a designed contextual encoder to semantically match distinct medical intents to achieve accurate intent recognition. Based on the recognition results, (iii) agent collaboration dynamically activates the corresponding paramedical agents. Then, a decision analysis module incorporates the outputs of multiple agents and generates integrated analysis results.
3.1 Query Rewriting
In this phase, we first process the initial query through the query input processor. The processor is based on a set of syntactic regularization algorithms that combine LLMs with a predefined ruleset $R$ (please refer to Appendix A.1 for details). The core concept is to check, optimize, and normalize user input to $Q_{std}$. Algorithm 1 shows the corresponding program. Then, a key element extractor distills critical information from $Q_{std}$, such as symptoms, condition descriptions, and medical histories, to form the element set $E_{i}$. In the Retrieval-Augmented Generation (RAG), we build an indexed database of 1,095 expert-reviewed medical guidelines and retrieve documents related to these elements using a semantic retrieval method  Lewis et al. (2021), forming the document set $D_{ref}$. After that, the recall analysis module selects the conforming documents by prompting guidance, which is fed into the LLM-based query prompter along with $Q_{std}$. The prompter optimizes the empirical information and efficient decomposition of composite intents. Eventually, we design a refined query constructor to merge and integrate the generated subqueries. It filters and reconstructs multiple subqueries $Q_{gen}$ with a predefined ruleset (please refer to Appendix A.2 for details) to ensure semantically complete and formally uniform output.
Algorithm 1 LLM-Based Input Standardization
  Output: Standardized query $Q_{std}$
  while not $converged$ do
     $converged\leftarrow\text{True}$
     for each rule $r\in\mathcal{R}$ do
        $Q_{new}\leftarrow\mathcal{M}(Q_{cur},r)$
        if $Q_{new}\neq Q_{cur}$ then
           $Q_{cur}\leftarrow Q_{new}$
           $converged\leftarrow\text{False}$
        end if
     end for
  end while
  return $Q_{std}\leftarrow Q_{cur}$
3.2 Intent Recognition
After the query rewriting, the optimized query $Q_{opt}$ is matched with a set of intent prototype embeddings $E_{i}$ generated by a contextual encoder, which is designed to capture the semantic features of different medical intents. The encoder is constructed on top of BioBERT Lee et al. (2019) to learn prototype representations by performing fine-grained intent classification. We add a fully connected layer after the embedding layer with output dimensions aligned to the 17 medical intent categories and generate the corresponding probability distributions via a softmax activation function. Specifically, the contextual encoder maps the optimized query together with the intent embeddings into a 768-dimensional embedding space. It computes the cosine similarity $S_{ij}$ between the query and each intent embedding $E_{i}$ with the following formula:
$\alpha_{ij}=\frac{\exp(S_{ij})}{\sum_{l=1}^{17}\exp(S_{il})}.$ (2)
3.3 Agent Collaboration
The corresponding agents are dynamically activated when the framework recognizes relevant and valid intents. Considering comprehensiveness and diversity, agent collaboration covers systematic healthcare services spanning pre-diagnosis, diagnosis, medicament, and post-diagnosis. Also, an LLM-based decision analysis module is employed for information integration and summarization.
Specifically speaking, the pre-diagnosis agent equips the patient repository, which is a relational database Codd (1970) for storing historical patient visit records. These records can be utilized for health assessments, identifying potential risks, and recommending appropriate departments to visit.
The diagnosis agent incorporates 506 high-quality cases provided in the medical record database Fan et al. (2024). We employ a hybrid retrieval scheme Sawarkar et al. (2024) to obtain demonstrations similar to the patient’s condition, thus providing reliable evidence regarding precise treatment recommendations in symptom analysis and etiological ranking intents. The hybrid retrieval consists of two main components: keyword retrieval and vector retrieval. First, the framework extracts a subset of documents $D_{\text{slice}}$ containing the query keyword from the document collection $D$ by the keyword retrieval of Eq. (1):
$D_{\text{slice}}=\{d\in D\mid\text{KeywordMatch}(Q,d)=True\}.$ (4)
Afterward, MedAide calculates the similarity $S(Q,d)$ between the query $Q$ and the document $d$ according to Eq. (2) by vector retrieval and retains the documents whose similarity is higher than a threshold set $\tau$, forming a subset $D_{\text{match}}$:
$S(Q,d)=\frac{E_{Q}\cdot E_{D}}{\|E_{Q}\|\|E_{D}\|},$ (5)
$D_{\text{match}}=\{d\in D\mid S(Q,d)>\tau\}.$ (6)
Subsequently, the keyword matching results in $D_{\text{slice}}$, and the vector retrieval results in $D_{\text{match}}$ are merged to obtain the final set of relevant documents $D_{\text{final}}$. These documents will provide empirical support for diagnostic recommendations:
$D_{\text{final}}=D_{\text{slice}}\cup D_{\text{match}}$ (7)
In this case, the diagnosis agent further interprets the ancillary examination results to help patients understand their health status. Meanwhile, it provides a second diagnostic opinion to minimize the risk of misdiagnosis.
The medicament agent utilizes 26,684 medication data extracted from the PubMed library Pud (2024) and provides patients with comprehensive medication counseling through keyword and semantic hybrid retrieval. The holistic counseling includes drug usage, dosage recommendations, potential drug interactions, and contraindications assessments to ensure the safety and efficacy of the medication. In addition, it helps patients understand potential side effects and suggests the best timing for medication administration to optimize treatment outcomes.
The post-diagnosis agent focuses on providing personalized rehabilitation advice and emotional support. It tracks and manages the recovery process by updating the patient repository with patient information.
The Decision Analysis Module (DAM) with the CoT properties combines the output of specific activated agents with the patient’s medical histories to dynamically adjust the treatment protocols and formulate integrated rehabilitation plans. The final answer generated by DAM based on $D_{\text{final}}$, the prompt, and the agent output $A_{i}$ is expressed as follows:
$Answer=\text{DAM}(A_{i},\text{prompt},D_{\text{final}}).$ (8)
Models BLEU-1 BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L GLEU
ZhongJing2/ + MedAide 9.06/9.18 2.68/2.81 17.54/19.83 1.70/2.86 10.04/11.74 3.66/4.19
Meditron-7B/ + MedAide 3.76/4.94 0.91/1.30 9.63/10.51 0.64/0.68 4.52/5.21 1.74/2.07
HuatuoGPT-II/ + MedAide 14.6/15.18 5.87/6.63 26.81/31.84 5.26/9.88 13.57/19.04 6.43/7.25
Baichuan4/ + MedAide 12.97/15.95 5.39/7.85 22.68/33.60 4.68/11.69 13.33/20.75 5.57/7.98
LLama-3.1-8B/ + MedAide 10.95/15.36 2.13/6.78 22.47/32.54 2.62/9.41 11.45/19.29 4.37/6.45
GPT-4o/ + MedAide 15.28/15.93 6.33/7.56 26.89/30.78 5.95/8.65 14.15/17.64 6.34/7.31
Table 1: Comparison results on the Pre-Diagnosis benchmark.
Models BLEU-1 BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L GLEU
ZhongJing2/ + MedAide 10.50/11.72 4.80/5.58 24.08/24.59 4.93/4.79 12.94/11.96 5.85/5.79
Table 2: Comparison results on the Diagnosis benchmark.
4 Experiments
4.1 Datasets and Implementation Details
All experiments are conducted in the zero-shot setting. To reflect the real-world demands for healthcare services, we conduct four different benchmarks across 17 types of medical intents, including Pre-Diagnosis, Diagnosis, Medicament, and Post-Diagnosis benchmarks. Each benchmark contains 500 composite intent instances. Please refer to Figures 7,  8,  9, and  10. in the Appendix for the specific intent taxonomy. For a comprehensive overview of the prompts used during the query rewriting and agent collaboration phases within MedAide, refer to Appendix  B. Additionally, to enable fine-grained medical intent recognition, we develop a contextual encoder training dataset consisting of 2,800 training samples, 300 validation samples, and 300 testing samples. Further configuration details can be found in Appendix A.3.
Models BLEU-1 BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L GLEU
ZhongJing2/ + MedAide 11.83/11.91 4.48/4.57 21.61/24.23 4.65/5.91 13.55/14.75 5.32/6.38
Table 3: Comparison results on the Medicament benchmark.
Models BLEU-1 BLEU-2 ROUGE-1 ROUGE-2 ROUGE-L GLEU
ZhongJing2/ + MedAide 12.97/19.91 4.86/6.86 25.31/31.98 4.21/4.48 12.44/15.00 5.63/7.22
LLama-3.1-8B/ + MedAide 21.85/22.95 7.42/7.68 35.36/39.71 5.22/7.97 15.02/17.93 7.65/8.01
Table 4: Comparison results on the Post-Diagnosis benchmark.
4.2 Model Zoo
We compare a series of state-of-the-art (SOTA) models for comprehensive evaluations. Among medical LLMs, ZhongJing2 Shi (2023) is a traditional Chinese medicine model based on Qwen-1.5-1.8B with complete training procedures. Meditron-7B Chen et al. (2023b) performs medically relevant continuous pre-training on Llama-2-7B to extend the breadth and depth of the model in medical knowledge. HuatuoGPT-II (13B) Chen et al. (2023a) employs a one-stage unified approach for domain adaptation to improve medical expertise and model applicability. In general-purpose models, Baichuan4 AI (2024) optimizes long texts by scaling law and combines reinforcement learning techniques to improve reasoning ability. Llama-3.1-8B Dubey et al. (2024) relies on the grouped query attention mechanism to enhance the inference efficiency and performs well on multilingual tasks. GPT-4o OpenAI (2024) shows excellent language comprehension and generation capabilities and excels in handling complex tasks.
4.3 Comparision with SOTA Methods
As a plug-and-play framework, we combine MedAide with the baseline models to provide comprehensive evaluations by different metrics, including BLEU-1/2 (%), ROUGE-1/2/L (%), and GLEU (%). Comparison results on the four medical benchmarks are shown below.
Results on Pre-Diagnosis Benchmark. As Table 1 shows, our framework consistently improves the performance across metrics for all models. (i) For instance, MedAide-based HuatuoGPT-II and Baichuan4 show significant relative improvements in ROUGE-L scores of 40.3% and 55.6%, respectively, suggesting multidimensional health risk assessments for patient cases. (ii) In addition, MedAide helps general LLama-3.1-8B to enhance the diversity and precision of diagnostic content generation, as reflected in the average 79.15% increase of BLEU-1/2 relative to the vanilla baseline. (iii) Despite the performance constraints of ZhongJing2 and Meditron-7B in tackling complex medical scenarios due to the scaling law, our framework still provides sustained performance gains.
Results on Diagnosis Benchmark. Table 2 shows the diagnosis task results containing six composite intent scenarios. (i) Combined with our framework, Meditron-7B and HuatuoGPT-II outperform vanilla baselines in response accuracy and completeness, as evidenced by significant improvements across metrics. These gains show that MedAide assists medical LLMs in giving more comprehensive symptom analyses and treatment suggestions with enhanced hierarchical interpretation capabilities regarding examination results. (ii) Meanwhile, the general models (e.g., LLama-3.1-8B and GPT-4o) equipped with MedAide enhance the specialization and adaptability in healthcare services. (iii) We observe slight performance drops (less than 1%) in ZhongJing-2 on ROUGE-2/L and GLEU. The potential reason is that the model has difficulty in the query rewriting phase to grasp the multifaceted knowledge of the retrieval-augmented generation due to the limited capacity, causing sub-optimal prevention guidelines and second opinions.
Pre-Diagnosis Diagnosis Medicament Post-Diagnosis
Components ROUGE-L GLEU ROUGE-L GLEU ROUGE-L GLEU ROUGE-L GLEU
Full MedAide 17.64 7.31 23.67 12.69 52.71 37.09 18.12 9.89
w/o Query Rewriting 16.02 6.44 18.35 10.37 32.06 32.45 16.48 8.79
w/ GPT-4o Recognition 16.58 6.73 22.98 12.34 44.72 34.10 17.56 9.02
w/o Decision Analysis 17.22 7.22 23.14 12.54 48.72 35.25 18.04 9.56
Table 5: Ablation study results on four benchmarks.“w/” and “w/o” are short for with and without, respectively.
Benchmarks MedAide w/o QR w/ GPT-4 R
Pre-Diagnosis 0.76 0.62 0.47
Diagnosis 0.83 0.61 0.49
Medicament 0.80 0.49 0.51
Post-Diagnosis 0.63 0.56 0.37
Intent Aggregation 0.86 0.60 0.60
Table 6: Comparison of F1 scores (%) across five benchmarks. “QR” means the Query Rewriting. “GPT-4o R” means the GPT-4o Recognition.
Results on Medicament Benchmark. Table 3 provides the exploration results of different models on composite medication intents. (i) Attributed to the drug knowledge injection and prototype-guided embedding in MedAide, the reliability of medication counseling responses is consistently improved across all baselines. (ii) In this case, Baichuan4 and GPT-4o demonstrate the most significant overall gains. For example, the GLEU scores separately increase 28.91% and 28.48%. This finding indicates that MedAide facilitates LLMs with large-window contexts to comprehend rich drug information and provide specialized dosage recommendations. (iii) In the zero-shot reasoning pattern, Meditron-7B’s ROUGE-1/2/L remarkably enhances from 12.29/2.12/6.31% to 34.84/29.43/32.57%, verifying that our decision-analysis module provides favorable factual evidence in drug contraindication understanding.
Results on Post-Diagnosis Benchmark. (i) In the post-diagnosis applications from Table 4, we observe that most MedAide-based models are superior to vanilla baselines by large margins. (ii) GPT-4o performs even better in generating personalized rehabilitation recommendations, with a relative improvement of 39.23% in the BLEU-1 score, providing hopefully more reliable rehabilitation guidelines for patients. (iii) Meanwhile, Baichuan4 improves on the ROUGE-L metric by 21.35%, revealing the MedAide-based version’s advantages in processing dynamic medical data and producing progress tracking reports.
4.4 Expert Doctor Evaluation
Expert evaluations play a key role in the practical applications of medical models. We invite 6 doctors (each paid \$300) to select the winners of responses generated by different models before and after the introduction of MedAide by majority voting rule. The response content is holistically evaluated by considering factual accuracy, recommendation practicality, and humanistic care. (i) As shown in Figure 2, all the MedAide-based models exhibit more win rates in different benchmark tests, indicating the effectiveness and applicability of the proposed framework. (ii) MedAide not only improves the healthcare specialization of medical LLMs with different sizes, but also significantly enhances the coping ability of general-purpose models when dealing with complicated medical tasks.
4.5 Ablation Study
We perform systematic ablation studies to explore the effects of different components in Table 5. GPT-4o is chosen as the baseline to combine MedAide.
Necessity of Query Rewriting. Firstly, we remove the query rewriting procedure to assess its impact on performance. (i) The observed significant declines in all metrics indicate that query rewriting is crucial for ensuring that the input information is accurately interpreted by downstream modules with respect to intents. (ii) Additionally, incorporating factual information as context enhances the model’s ability to accurately comprehend user queries, mitigating hallucinations and reducing the generation of ambiguous interpretations.
Importance of Intent Recognition. We replace the learning-based contextual encoder with the prompt-based GPT-4o to explore the impact of different intent recognition strategies. (i) Despite the improvement over the vanilla baseline, GPT-4o recognition is a sub-optimal solution. We argue that the encoder trained with explicitly supervised information can align more purposefully to actual medical intents, producing more personalized and representative judgments than prompt engineering. (ii) Also, our default strategy has better flexibility and extensibility, which can be dynamically optimized according to specific healthcare scenarios.
Effectiveness of Decision Analysis. Furthermore, we directly assemble the different outputs produced by the activated agents to serve as an alternative candidate. The results from the bottom of Table 5 show consistent performance drops of the model on different medical intent understandings, proving the effectiveness of our method. A plausible explanation is that the decision analysis module not only summarizes the different agent outputs in an organized manner, but also provides comprehensive and accurate conclusions based on medical guidelines and patient histories.
Benchmarks MedAide (Ours) MedAgent
Pre-Diagnosis 15.93/17.64 14.33/15.26
Diagnosis 27.80/23.67 25.46/19.54
Medicament 45.13/52.17 28.69/36.07
Post-Diagnosis 26.83/18.12 25.45/16.69
Table 7: Comparison of BLEU-1/ROUGE-L scores (%) between MedAide and MedAgent across benchmarks.
4.6 Intent Detection Analysis
To further observe the intent detection performance, Table 6 compares the F1 scores computed between the intents recognized by MedAide and ground truths under different strategies. We randomly sample 100 instances from each benchmark to compose a more challenging Intent Aggregation benchmark. The core observations are as follows. (i) The complete framework achieves the best results, reaching even 86% performance on the Aggregation benchmark, exhibiting omni intent semantic understanding capabilities. (ii) Significant performance deteriorations occur at “w/o QR” and “w/ GPT-4o R”, implying the effectiveness of fine-grained decomposition of medical queries and the development of our tailored intent recognition mechanism.
Figure 4: The effect of different agent numbers on performance across the four medical benchmarks.
4.7 Collaboration Framework Comparison
Here, we compare the reproducible multi-agent framework MedAgent Tang et al. (2024). This medical framework mainly sets up four department-specific experts, including pediatrics, cardiology, pulmonology, and neonatology. Each expert is based on the GPT-4o implementation and works together to collaboratively generate a unified report as the final result. (i) From Table 7, our framework outperforms MedAgent on all four benchmarks. For instance, in the Medicament benchmark, MedAide beats MedAgent by large margins with absolute gains of 16.44/16.70% on the BLEU-1/ROUGE-L scores. This strength comes from our specialized medication retrieval, which provides a refined context of drug information for the model. (ii) Also, the benefits on Pre-Diagnosis and Post-Diagnosis tasks reflect that the proposed MedAide captures the key characteristics of diagnostic and nursing demands more efficiently.
4.8 Response Visualization Analysis
To intuitively compare the response quality of the different models, we show in Figure 3 the generated results of MedAide and SOTA methods under the same user query. HuatuoGPT-II provides extensive medical knowledge but lacks targeted recommendations. Although the MedAgent framework gives detailed treatment solutions through multidisciplinary perspectives, it falls short of coherent diagnostic reasoning. In contrast, our framework interprets patient outcomes in detail and provides personalized diagnoses and subsequent treatments. MedAide shows favorable usefulness in electrolyte imbalance and cardiovascular risk management.
4.9 Effect of Agent Number
Figure 4 illustrates the variation trend of model performance when the number of agents in the proposed MedAide changes. When the agent number is increased from 2 to 4, there are progressive improvements in the results on all benchmarks, suggesting that multi-agent collaboration helps to improve the model’s performance in multi-tasking scenarios. Conversely, when the agent number exceeds 4 does not bring significant gains due to introducing too much information redundancy.
5 Conclusion and Future Work
This paper introduces MedAide, a multi-agent framework for complex medical scenarios. Leveraging query rewriting, intent recognition, and agent collaboration, MedAide enhances the model’s comprehension of medical intents and demonstrates effectiveness across multiple benchmarks.
Future Work. We plan to incorporate complementary modalities (e.g., medical imaging) in the future to improve the potential of MedAide in multimodal diagnostics and applications.
Limitations
The MedAide framework, while making significant strides in integrating large-scale medical agents with real clinical environments, has its limitations. Currently, the framework incorporates 26,684 drug samples and 506 genuine clinical case records. Despite this scope, it lacks comprehensive coverage, particularly in research on rare diseases and specialized medications. Furthermore, as the initial versions primarily emphasized linguistic processing, future research aims to extend into multimodal capabilities, particularly the integration of medical imaging data, to enhance the framework’s ability to process visual information in clinical diagnostics. The reliance on OpenAI’s API may pose potential operational challenges, suggesting that future studies should explore using more efficient open-source models as viable alternatives. Despite these limitations, the MedAide framework provides a solid foundation for advancing the integration of artificial intelligence in clinical diagnostics.
Ethics Consideration
Ethical considerations are paramount in the application of medical agents to real-world clinical settings. We are fully aware of the potential impacts of our research and have taken deliberate actions to address these issues. To enhance transparency, we are committed to publicly making the drug data and medical records used in our study accessible. This will enable other researchers to validate our findings and build upon our work, fostering collaboration and advancement in this field.
We are acutely aware of the necessity for privacy and data protection. All data utilized has undergone thorough de-identification, with all sensitive information removed, and verified by a partnering medical institution. We invite doctors to perform only evaluations of model responses without involving any form of human subject research. All participants are compensated \$300 for their work, which strictly adheres to the minimum hourly rate for the region in which the work is performed. For the utilization of healthcare-related data, we strictly follow the license agreements of publicly available databases. For the constructed data, we have undergone an internal ethical review by the ethics review board of our partnering medical institutions and are licensed and approved.
Acknowledgment
This work is supported in part by the National Key R&D Program of China under Grant 2021ZD0113502, in part by the Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0103, and in part by the Shanghai Municipal Science and Technology Committee of Shanghai Outstanding Academic Leaders Plan (No. 21XD1430300).
Appendix A Implementation Details
A.1 Query Normalization and Parsing
@@
  Let $Q=\{w_{1},w_{2},\dots,w_{n}\}$ be the sequence of words (tokens) in the query.
@@
Algorithm 2 first disambiguates the input query and divides the query into individual words or tokens. Next, the algorithm initializes an empty syntax tree and matches substrings in the query by Context-Free Grammar (CFG) rules, adding the conforming rules to the syntax tree. The algorithm keeps applying the rules recursively until the syntax tree covers the entire query and finally outputs the complete syntax tree.
A.2 Refined Query Construction
In the refined query constructor, we implement a set of rules (as shown in Figure 5) to ensure that the generated subqueries are of high quality both semantically and formally. These rules encompass subquery filtering, semantic overlap removal, query consolidation, grammatical normalization, intent prioritization, and format standardization. Together, these processes streamline the output, optimize the query structure, and prioritize critical medical intents, ensuring the final query is both semantically coherent and structurally consistent.

### v3

@@ 3 Methodology
Figure 2 illustrates the overall architecture of the proposed MedAide framework. The workflow comprises three key components: Regularization-guided Information Extraction (RIE) module, dynamic Intent Prototype Matching (IPM) module, and Rotation Agent Collaboration (RAC) mechanism.
@@
  Let $Q={w_{1},w_{2},\dots,w_{n}}$ be the sequence of words (tokens) in the query.
@@
Algorithm 2 LLM-Based Input Standardization
  Output: Standardized query $Q_{std}$
  while not $converged$ do
     $converged\leftarrow\text{True}$
     for each rule $r\in\mathcal{R}$ do
        $Q_{new}\leftarrow\mathcal{M}(Q_{cur},r)$
        if $Q_{new}\neq Q_{cur}$ then
           $Q_{cur}\leftarrow Q_{new}$
           $converged\leftarrow\text{False}$
        end if
     end for
  end while
  return $Q_{std}\leftarrow Q_{cur}$
3.1 Regularization-guided Information Extraction
Due to the inherent complexity and ambiguity of natural language expressions in medical contexts, the precise identification of complex medical intents and information extraction presents significant challenges [33]. Previous works have utilized rule-based approaches [5] and traditional natural language processing techniques to achieve certain levels of medical text understanding, but they suffer from the following limitations: (1) inadequate handling of composite medical intents that contain multiple interconnected clinical concerns; (2) insufficient semantic understanding of domain-specific medical terminology and contextual relationships. To address these challenges, we propose a Regularization-guided Information Extraction (RIE) module that leverages syntactic regularization algorithms combined with large language models to achieve robust medical query understanding and intent decomposition.
In this phase, we first process the initial query through the query input processor. The processor is based on a set of syntactic regularization algorithms that combine LLMs with a predefined ruleset. In this case, the core concept of RIE is to check, optimize, and normalize user input to $Q_{std}$. Algorithm 1 shows the corresponding program. First, the algorithm disambiguates the input query and divides the query into individual words or tokens. Next, the algorithm initializes an empty syntax tree and matches substrings in the query by context-free grammar rules, adding the conforming rules to the syntax tree. The algorithm keeps applying the rules recursively until the syntax tree covers the entire query and finally outputs the complete syntax tree.
Then, a key element extractor distils critical information from $Q_{std}$, such as symptoms, condition descriptions, and medical histories, to form the element set $E_{i}$. In the Retrieval-Augmented Generation (RAG), we build an indexed database of 1,095 expert-reviewed medical guidelines and retrieve documents related to these elements using a semantic retrieval method  [21], forming the document set $D_{ref}$. Subsequently, the recall analysis module identifies relevant documents through prompt-guided selection based on BERT-Score [44], which is fed into the LLM-based query prompter along with $Q_{std}$. The prompter optimizes the empirical information and efficient decomposition of composite intents. Eventually, we design a refined query constructor to merge and integrate the generated subqueries. To ensure that the generated subqueries maintain high quality in both semantic meaning and formal structure, we implement a set of rules (as shown in Figure 3) in the refined query constructor using Algorithm 2. These rules encompass subquery filtering, semantic overlap removal, query consolidation, grammatical normalization, intent prioritization, and format standardization. Specifically, the constructor filters and reconstructs multiple subqueries $Q_{gen}$ with these predefined rulesets to ensure semantically complete and formally uniform output. These processes streamline the output, optimize the query structure, and prioritize critical medical intents, ensuring that the final query is semantically coherent and structurally consistent.